# EBPO: Empirical Bayes Shrinkage for Stabilizing Group-Relative Policy Optimization

### v1

@@ 1 Introduction
Post-training for Large Language Models (LLMs) has increasingly shifted toward reinforcing reasoning using verifiable signals. Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a stable and reproducible paradigm for tasks such as mathematical reasoning and code generation, where objective correctness can serve as a ground-truth reward signal shao2024deepseekmath; guo2025deepseek. To avoid the computational overhead of training separate value networks, recent advancements have coalesced around group-based methods, most notably Group Relative Policy Optimization (GRPO) shao2024deepseekmath. By normalizing rewards within a sampled group of outputs for a given prompt, GRPO provides a computationally efficient baseline that has demonstrated significant improvements in model reliability shao2024deepseekmath; yu2025dapo.
@@
Despite its success, GRPO faces inherent limitations regarding the stability of its advantage estimator. Because the baseline is derived solely from the local group mean, it is susceptible to high variance when the group size ($G$) is small yu2025dapo. Furthermore, GRPO lacks robustness in “saturated” regimes: if a model fails all attempts for a difficult prompt (all rewards are zero), the relative advantage vanishes, resulting in a null gradient and a wasted training step liu2025understanding. While recent approaches like DAPO yu2025dapo attempt to mitigate this volatility by filtering or reweighting unstable updates, they do so by effectively discarding the problematic examples, leading to unavoidable data waste. Consequently, standard GRPO often necessitates larger group sizes to suppress noise without losing data, thereby drastically increasing the computational cost of training shao2024deepseekmath.
@@
In this work, we introduce Empirical Bayes Policy Optimization (EBPO), which reframes advantage estimation through the lens of Empirical Bayes (EB) inference robbins1992empirical. Classical EB theory suggests that when estimating parameters for parallel tasks, one can reduce estimation error by assuming those parameters share a common underlying distribution robbins1992empirical. Applying this to RLVR, we postulate that the latent success probability of a prompt is drawn from a global distribution characterized by the policy’s historical performance.
@@
EBPO replaces the purely local GRPO baseline with a shrinkage estimator that pulls the noisy group mean toward a global mean, $\mu_{glob}$. The degree of this shrinkage is determined dynamically by the ratio of within-group variance to between-group variance. This formulation allows EBPO to distinguish between failing a globally “hard” task (consistent with the prior) versus failing an “easy” task (deviating from the prior), assigning adaptive penalty signals accordingly. To ensure scalability, we estimate these global priors dynamically using Welford’s online algorithm welford1962note.
@@
We validate EBPO through both theoretical analysis and extensive empirical evaluation. Theoretically, we prove that the EBPO baseline achieves a strictly lower Mean Squared Error (MSE) than the standard sample mean used in GRPO and provides informative, non-zero gradients even when group rewards are saturated. Empirically, we evaluate EBPO on various datasets with different scales of base models. Our results show that EBPO consistently outperforms baselines on challenging benchmarks such as AIME-2024 li2024numinamath and OlympiadBench he2024olympiadbench. Furthermore, we demonstrate that EBPO is highly sample-efficient, outperforming GRPO by over 11% in extremely resource-constrained settings ($G=8$), while effectively leveraging curriculum learning via difficulty-stratified sampling.
@@ 2 Related Work
Previous work has shown that RLVR is effective in enhancing reasoning and factual correctness in LLMs shao2024deepseekmath; guo2025deepseek. By leveraging automatically checkable signals, RLVR enables stable and reproducible optimization for reasoning tasks. A variety of algorithmic variants were proposed, including GRPO-based approaches that integrate verifiable rewards into group-based policy optimization frameworks shao2024deepseekmath; liu2025understanding; yu2025dapo; cui2025entropy; zheng2025group; zhou2025mixture. These methods demonstrate that verifiable reward signals can significantly improve model reliability.
Among these GRPO variants, a broad range of studies focus on improving the effectiveness and efficiency of RLVR training, such as by proposing novel learning objectives liu2025understanding; zhou2025disco; yu2025dapo; yue2025vapo; liu2025uniform, introducing entropy-based mechanisms to encourage broader exploration cui2025entropy; wang2025beyond; zhang2025edge; le2025no; yang2025let, or designing more robust reward functions jia2025autorubric; li2025semantically; xiong2026token. However, few have considered the estimation of the mean or standard deviation of rewards during advantage computation, which can substantially influence group-level differences. Our work addresses this gap by employing an Empirical Bayes method to more robustly estimate the mean and standard deviation of group rewards, thereby stabilizing the training dynamics of GRPO.
@@
Empirical Bayes provides a framework to handle estimation in the face of parallel, noisy data. It is a powerful methodology that blends the data-driven nature of frequentist statistics with the hierarchical modeling structure of Bayesian inference, offering a principled approach to regularization and robust estimation. Instead of specifying the prior hyperparameters in advance, Empirical Bayes uses the observed data to estimate them robbins1992empirical. Later, statisticians found that shrinkage estimators like the James-Stein estimator james1961estimation can be viewed as Empirical Bayes estimators with a Gaussian prior. Empirical Bayes methods have been widely adopted for hyperparameter tuning in Bayesian reinforcement learning reisinger2008online and prior estimation in multi-task meta-learning hu2020empirical; grant2018recasting. Additionally, these methods facilitate uncertainty estimation in deep neural networks krishnan2020specifying and the development of Bayesian dialogue agents through data-driven parameter initialization lee-etal-2023-empirical.
@@ 3.1 Preliminaries
RLVR provides a memory-efficient framework for optimizing reasoning policies $\pi_{\theta}$ in domains where the correctness of a response $o$ can be objectively determined by a verifier shao2024deepseekmath. Unlike traditional actor-critic methods that require a learned value network, RLVR utilizes group-relative advantage estimation to provide a baseline for policy updates.
@@
In classical Bayesian inference, a prior distribution is specified before any data is observed. In Empirical Bayes (EB), the prior is instead estimated from the data itself robbins1992empirical. Consider the general problem of simultaneously estimating a set of $M$ related parameters $\theta_{1},\dots,\theta_{M}$ based on independent observations $y_{1},\dots,y_{M}$, where each $y_{m}$ serves as a noisy estimate of $\theta_{m}$. Rather than estimating each parameter in isolation—which often yields high variance when data is sparse—EB assumes that the parameters are exchangeable and drawn from a common global prior distribution: $\theta_{m}\sim\mathcal{N}(\mu,\tau^{2})$, where $\mu$ is the global mean and $\tau^{2}$ is the between-parameter variance (or prior variance). By estimating these global hyperparameters $(\mu,\tau^{2})$ from the marginal distribution of the entire dataset, we can construct a posterior estimate for each individual $\theta_{m}$. This results in a shrinkage estimator, which “pulls” individual, noisy observations toward the global mean $\mu$.
@@ 3.2 Empirical Bayes Policy Optimization (EBPO)
A primary challenge in post-training LLMs on reasoning tasks, such as mathematics or formal logic, is the sparsity and high variance of rewards uesato2022solving; zelikman2022star; lightman2023lets; shao2024deepseekmath; yu2025dapo. In GRPO methods, the advantage is computed by normalizing rewards within a group of completions for a single prompt. However, when a prompt is excessively difficult, the model may fail all attempts (i.e., all rewards are zero), resulting in a null gradient. Conversely, for trivial prompts, the model might succeed in all attempts, again providing no relative signal for improvement.
@@
To ensure training stability and scalability, we maintain running estimates of $\mu_{glob}$, $\sigma^{2}$, and $\tau^{2}$ using Welford’s Online Algorithm welford1962note. This avoids the noise associated with small-batch statistics. For each step, the advantage for a completion $(q_{m},o_{m,i},r_{m,i})$ is computed as the deviation from this regularized baseline:
@@ 4.1 Experimental Setup
We use the DAPO-Math-17K dataset as the training corpusyu2025dapo. For evaluation, we select a suite of competitive math reasoning benchmarks, including AIME2024 li2024numinamath, AIME2025 li2024numinamath, AMC23 li2024numinamath, Math-500 hendrycks2021measuring, and OlympiadBench he2024olympiadbench. Each model is evaluated using the Pass@1 metric. To ensure statistical robustness, we repeat each evaluation set 32 times with different random seeds and report the average Pass@1 across runs.
@@
We adopt a diverse set of LLMs with different architectures and parameter scales. Specifically, we use LLaMA3.1-8B dubey2024llama, Qwen3-8B, and Qwen3-14B yang2025qwen3. This selection spans multiple LLM families (LLaMA and Qwen) and model sizes, providing a robust and comprehensive testbed for evaluating our agent workflow and training recipe.
@@
We compare our proposed EBPO method with several representative baselines, including Naive GRPO shao2024deepseekmath, DAPO yu2025dapo, Dr. GRPO liu2025understanding, and EntropyMech cui2025entropy. For a fair comparison, all methods are trained under identical conditions: the same training data ordering, batch size, and optimization configuration are applied across all baselines and our EBPO method. Specific configuration details for each baseline are provided in Appendix 6.
@@
To implement the clustered sampling optimization in Section 3.3.3, we evaluate two sampling strategies: clustering by topic and clustering by difficulty, which correspond to EBPO-topic and EBPO-diff, respectively. For difficulty-based clustering, we use the base model’s pass rate as the difficulty metric, grouping problems of similar success rates together. For topic-based clustering, we employ GPT-4.1 achiam2023gpt to annotate each problem with one topic label. Details are provided in Appendix 6.

### v2

@@ 1 Introduction
Post-training for Large Language Models (LLMs) has increasingly shifted toward reinforcing reasoning using verifiable signals. Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a stable and reproducible paradigm for tasks such as mathematical reasoning and code generation, where objective correctness can serve as a ground-truth reward signal (shao2024deepseekmath; guo2025deepseek). To avoid the computational overhead of training separate value networks, recent advancements have coalesced around group-based methods, most notably Group Relative Policy Optimization (GRPO) (shao2024deepseekmath). By normalizing rewards within a sampled group of outputs for a given prompt, GRPO provides a computationally efficient baseline that has demonstrated significant improvements in model reliability (shao2024deepseekmath; yu2025dapo).
@@
Despite its success, GRPO faces inherent limitations regarding the stability of its advantage estimator. Because the baseline is derived solely from the local group mean, it is susceptible to high variance when the group size ($G$) is small (yu2025dapo). Furthermore, GRPO lacks robustness in “saturated” regimes: if a model fails all attempts for a difficult prompt (all rewards are zero), the relative advantage vanishes, resulting in a null gradient and a wasted training step (liu2025understanding). While recent approaches like DAPO (yu2025dapo) attempt to mitigate this volatility by filtering or reweighting unstable updates, they do so by effectively discarding the problematic examples, leading to unavoidable data waste. Consequently, standard GRPO often necessitates larger group sizes to suppress noise without losing data, thereby drastically increasing the computational cost of training (shao2024deepseekmath).
@@
In this work, we introduce Empirical Bayes Policy Optimization (EBPO), which reframes advantage estimation through the lens of Empirical Bayes (EB) inference (robbins1992empirical). Classical EB theory suggests that when estimating parameters for parallel tasks, one can reduce estimation error by assuming those parameters share a common underlying distribution (robbins1992empirical). Applying this to RLVR, we postulate that the latent success probability of a prompt is drawn from a global distribution characterized by the policy’s historical performance.
@@
EBPO replaces the purely local GRPO baseline with a shrinkage estimator that pulls the noisy group mean toward a global mean, $\mu_{glob}$. The degree of this shrinkage is determined dynamically by the ratio of within-group variance to between-group variance. This formulation allows EBPO to distinguish between failing a globally “hard” task (consistent with the prior) versus failing an “easy” task (deviating from the prior), assigning adaptive penalty signals accordingly. To ensure scalability, we estimate these global priors dynamically using Welford’s online algorithm (welford1962note).
@@
We validate EBPO through both theoretical analysis and extensive empirical evaluation. Theoretically, we prove that the EBPO baseline achieves a strictly lower Mean Squared Error (MSE) than the standard sample mean used in GRPO and provides informative, non-zero gradients even when group rewards are saturated. Empirically, we evaluate EBPO on various datasets with different scales of base models. Our results show that EBPO consistently outperforms baselines on challenging benchmarks such as AIME-2024 (li2024numinamath) and OlympiadBench (he2024olympiadbench). Furthermore, we demonstrate that EBPO is highly sample-efficient, outperforming GRPO by over 11% in extremely resource-constrained settings ($G=8$), while effectively leveraging curriculum learning via difficulty-stratified sampling.
@@ 2 Related Work
Previous work has shown that RLVR is effective in enhancing reasoning and factual correctness in LLMs (shao2024deepseekmath; guo2025deepseek). By leveraging automatically checkable signals, RLVR enables stable and reproducible optimization for reasoning tasks. A variety of algorithmic variants were proposed, including GRPO-based approaches that integrate verifiable rewards into group-based policy optimization frameworks (shao2024deepseekmath; liu2025understanding; yu2025dapo; cui2025entropy; zheng2025group; zhou2025mixture). These methods demonstrate that verifiable reward signals can significantly improve model reliability.
Among these GRPO variants, a broad range of studies focus on improving the effectiveness and efficiency of RLVR training, such as by proposing novel learning objectives (liu2025understanding; zhou2025disco; yu2025dapo; yue2025vapo; liu2025uniform), introducing entropy-based mechanisms to encourage broader exploration (cui2025entropy; wang2025beyond; zhang2025edge; le2025no; yang2025let), or designing more robust reward functions (jia2025autorubric; li2025semantically; xiong2026token). However, few have considered the estimation of the mean or standard deviation of rewards during advantage computation, which can substantially influence group-level differences. Our work addresses this gap by employing an Empirical Bayes method to more robustly estimate the mean and standard deviation of group rewards, thereby stabilizing the training dynamics of GRPO.
@@
Empirical Bayes provides a framework to handle estimation in the face of parallel, noisy data. It is a powerful methodology that blends the data-driven nature of frequentist statistics with the hierarchical modeling structure of Bayesian inference, offering a principled approach to regularization and robust estimation. Instead of specifying the prior hyperparameters in advance, Empirical Bayes uses the observed data to estimate them (robbins1992empirical). Later, statisticians found that shrinkage estimators like the James-Stein estimator (james1961estimation) can be viewed as Empirical Bayes estimators with a Gaussian prior. Empirical Bayes methods have been widely adopted for hyperparameter tuning in Bayesian reinforcement learning (reisinger2008online) and prior estimation in multi-task meta-learning (hu2020empirical; grant2018recasting). Additionally, these methods facilitate uncertainty estimation in deep neural networks (krishnan2020specifying) and the development of Bayesian dialogue agents through data-driven parameter initialization (lee-etal-2023-empirical).
@@ 3.1 Preliminaries
RLVR provides a memory-efficient framework for optimizing reasoning policies $\pi_{\theta}$ in domains where the correctness of a response $o$ can be objectively determined by a verifier (shao2024deepseekmath). Unlike traditional actor-critic methods that require a learned value network, RLVR utilizes group-relative advantage estimation to provide a baseline for policy updates.
@@
In classical Bayesian inference, a prior distribution is specified before any data is observed. In Empirical Bayes (EB), the prior is instead estimated from the data itself (robbins1992empirical). Consider the general problem of simultaneously estimating a set of $M$ related parameters $\theta_{1},\dots,\theta_{M}$ based on independent observations $y_{1},\dots,y_{M}$, where each $y_{m}$ serves as a noisy estimate of $\theta_{m}$. Rather than estimating each parameter in isolation—which often yields high variance when data is sparse—EB assumes that the parameters are exchangeable and drawn from a common global prior distribution: $\theta_{m}\sim\mathcal{N}(\mu,\tau^{2})$, where $\mu$ is the global mean and $\tau^{2}$ is the between-parameter variance (or prior variance). By estimating these global hyperparameters $(\mu,\tau^{2})$ from the marginal distribution of the entire dataset, we can construct a posterior estimate for each individual $\theta_{m}$. This results in a shrinkage estimator, which “pulls” individual, noisy observations toward the global mean $\mu$.
@@ 3.2 Empirical Bayes Policy Optimization (EBPO)
A primary challenge in post-training LLMs on reasoning tasks, such as mathematics or formal logic, is the sparsity and high variance of rewards (uesato2022solving; zelikman2022star; lightman2023lets; shao2024deepseekmath; yu2025dapo). In GRPO methods, the advantage is computed by normalizing rewards within a group of completions for a single prompt. However, when a prompt is excessively difficult, the model may fail all attempts (i.e., all rewards are zero), resulting in a null gradient. Conversely, for trivial prompts, the model might succeed in all attempts, again providing no relative signal for improvement.
@@
To ensure training stability and scalability, we maintain running estimates of $\mu_{glob}$, $\sigma^{2}$, and $\tau^{2}$ using Welford’s Online Algorithm (welford1962note). This avoids the noise associated with small-batch statistics. For each step, the advantage for a completion $(q_{m},o_{m,i},r_{m,i})$ is computed as the deviation from this regularized baseline:
@@ 4.1 Experimental Setup
We use the DAPO-Math-17K dataset as the training corpus (yu2025dapo). For evaluation, we select a suite of competitive math reasoning benchmarks, including AIME2024 (li2024numinamath), AIME2025 (li2024numinamath), AMC23 (li2024numinamath), Math-500 (hendrycks2021measuring), and OlympiadBench (he2024olympiadbench). Each model is evaluated using the Pass@1 metric. To ensure statistical robustness, we repeat each evaluation set 32 times with different random seeds and report the average Pass@1 across runs.
@@
We adopt a diverse set of LLMs with different architectures and parameter scales. Specifically, we use LLaMA3.1-8B (dubey2024llama), Qwen3-8B, and Qwen3-14B (yang2025qwen3). This selection spans multiple LLM families (LLaMA and Qwen) and model sizes, providing a robust and comprehensive testbed for evaluating our agent workflow and training recipe.
@@
We compare our proposed EBPO method with several representative baselines, including Naive GRPO (shao2024deepseekmath), DAPO (yu2025dapo), Dr. GRPO (liu2025understanding), and EntropyMech (cui2025entropy). For a fair comparison, all methods are trained under identical conditions: the same training data ordering, batch size, and optimization configuration are applied across all baselines and our EBPO method. Specific configuration details for each baseline are provided in Appendix 6.
@@
To implement the clustered sampling optimization in Section 3.3.3, we evaluate two sampling strategies: clustering by topic and clustering by difficulty, which correspond to EBPO-topic and EBPO-diff, respectively. For difficulty-based clustering, we use the base model’s pass rate as the difficulty metric, grouping problems of similar success rates together. For topic-based clustering, we employ GPT-4.1 (achiam2023gpt) to annotate each problem with one topic label. Details are provided in Appendix 6.